OpenStudy (anonymous):

Ten years hence Rajesh’s age would be 15 years more than twice his friend Sanjay’s age. 10 years ago Rajesh’s age was 35 years more than twice his friend’s age. After how many years would Rajesh’s age be exactly twice of his friend’s age?

OpenStudy (anonymous):

After 25years...

OpenStudy (anonymous):

could you plz explain

OpenStudy (anonymous):

r-rajesh's age as of now; s-sanjay's from the first clause of the qn(after 10 yrs): (r+10)=15+2(s+10) so now r=2s+25 (rearranging) in fact both the clauses will give the same eqn. when s=1; r=27 after 25 years: s=26; r=52 (which is double)

OpenStudy (anonymous):

k thanks